OpenStudy (anonymous):

*radical*cos^2(*theta*/2)

OpenStudy (anonymous):

\[\sqrt{\cos^{2}(\theta \div2) }\]

OpenStudy (anonymous):

Rewrite the expression in nonradical form without using absolute values for the indicated values of \[\theta\] \[0<\]\[\theta <\pi\]